# Can Large Language Models be Good Path Planners? A Benchmark and Investigation on Spatial-temporal Reasoning

### v2

@@ 1 Introduction
Our experimental results suggest that the GPT-4 model does possess some capability to perform spatial reasoning, when it is prompted effectively. Specifically, the action-and-effect prompting leads to 21.5% improvement in success rate than the naive few-shot prompting, implying that guiding LLMs to reason about the situated knowledge can help them make better immediate decisions. Prompting GPT-4 to reason step-by-step (i.e., CoT) further enhances its performance by 3%, whereas additionally allowing the model to perceive the environment signals (i.e., ReAct) yields the best success rate, 96.1%. However, we also note that this success is mostly limited to relatively “local” reasoning; for example, the ReAct-prompted GPT-4 is shown to plan only a few steps ahead and have to frequently adjust its plan based on the environment signals. In summary, this indicates that even the state-of-the-art GPT-4 still falls short in long-term, temporal reasoning. In contrast, while the fine-tuned LLMs can achieve a close-to-perfect success rate when planning in environments similar to where they were trained (i.e., in-distribution environments), they suffer from the lack of generalization and cannot perform well in environments with different grid sizes or different number of obstacles. We will release our benchmark, along with the source code to controllably synthesize the grid environment, as well as our LLM implementation, to the public for future advancement  [^1] Our source code and data are available on the following link..
@@ 2.2 Data Generation Process
(Patel & Pavlick, 2022)
Understanding
None
2D Grid
Synthetic
No
Yes
Ruis et al. (2020)
Reasoning
Short-horizon Planning
2D Grid
Synthetic
No
No
Shridhar et al. (2020)
Reasoning
Short-horizon planning
Embodied Environment
Human-annotated
No
No
Côté et al. (2018)
Reasoning
Short-horizon planning
Embodied Environment
Synthetic
No
Yes
Shridhar et al. (2021)
Reasoning
Short-horizon planning
Embodied Environment
Human-annotated
No
Yes
Valmeekam et al. (2022)
Reasoning
Short-horizon planning
Blocksworld
Synthetic
Yes
Yes
PPNL (ours)
Reasoning
Long-horizon Planning
2D Grid
Synthetic
Yes
Yes
@@ 3 Methodologies
ReAct Prompting: Finally, we emulate how an agent can interleavedly perceive environment signals, reason about its plan, and take an action, following the approach of ReAct (Yao et al., 2023). Specifically, we provide two types of signals to the LLM: (1) informing the LLM when it has successfully reached the goal, and (2) reminding the LLM when it is one step ahead of any obstacle, as well as providing its current location. Particularly for the latter type, it allows the LLM to prevent potential failures and correct its ineffective plan. It is important to note that both signals are “local”; that is, the agent will only observe its surroundings and will not be provided with a global view. Therefore, to succeed in the path planning task, it is still required to perform long-term spatial-temporal reasoning. Considering the prohibitive cost of prompting an LLM iteratively, in experiments, we allow a ReAct LLM to incorporate environment signals and adjust its plan for at most three trials.
@@ 4.1 In-Distribution Experiments
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@ 6-11 Obstacles
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@ 4.2 Out-of-Distribution Experiments
Fine-tuned models can generalize to simpler grids but not to more complex ones. We observed that the fine-tuned LLMs were able to generalize well to the 5x5 environments. We hypothesize that this is because 5x5 environments are inherently part of the 6x6 environments encountered in training. For the 7x7 environments, however, all of the fine-tuned models failed to perform well. This highlights OOD generalization as an important limitation of fine-tuned models.
@@ 5 Multi-Goal Path Planning Results
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@ 6-11 Obstacles
As far as OOD generalization is concerned, we notice a trend similar to that of the single-goal OOD generalization. T5-base is still able to perform well on the 5x5 environments; however, it failed catasrophically when faced with larger grids, while also experiencing slighter performance drops on the 6-11 obstacles set.
@@ 7 Related Work
Path and Motion Planning: Many works have assessed whether LLMs can serve as planners (Song et al., 2022; Huang et al., 2022a; Valmeekam et al., 2022). However, many arrive to the conclusion LLMs are unsuitable to be end-to-end planners (Liu et al., 2023; Chen et al., 2023; Xie et al., 2023; Silver et al., 2022). We perform an in-depth analysis to verify this claim. Robot motion and path planning is an exciting application for LLMs that has been explored recently (Ahn et al., 2022; Ding et al., 2023; Chen et al., 2023; Driess et al., 2023; Huang et al., 2022b). Nevertheless, applications of LLMs in the realm of robotics remain limited. Our benchmark can, hence, be particularly useful to the robotic path and motion planning community, as we provide a controlled environment for studying the capability of LLMs to serve as path planners, and their ability to generalize to novel situations, a skill that is crucial in a large number robotic applications.
@@ 8 Conclusions and Future Work
Acknowledgments
This project was supported by resources provided by the Office of Research Computing at George Mason University (URL https://orc.gmu.edu) and funded in part by grants from the National Science Foundation (Awards Number 1625039 and 2018631).
@@ A.1 Samples Generation
World Description
You are in a $\{N\}$ by $\{N\}$ world.
@@
Enumerating Goals
$p\{i\}$ is located at $\{(x_{i},y_{i})\}$
Single goal
Go from $\{(x_{init},y_{init})\}$ to $\{(x_{goal},y_{goal})\}$.
Multi Goal
Visit the following locations: $\{locations\}$.
Constrained Ordering
Visit $\{locations_{a}\}$ before $\{locations_{b}\}$
Initial Location
You are at $\{(x_{init},y_{init})\}$
@@ Appendix C Fine-tuned Models’ Performance on Sampled Test Sets
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@ 6-11 Obstacles
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@ Appendix D Additional Analysis
Success (normal-$\uparrow$) Optimal (normal-$\uparrow$) Exact Match (normal-$\uparrow$) Feasible (normal-$\uparrow$) Distance (normal-$\downarrow$) Unreachable Acc (normal-$\uparrow$)
@@
Naive few-shot (5)
418.85
3.91
0.012
Naive few-shot (10)
722.85
3.86
0.022
Naive few-shot (15)
1013.86
3.95
0.031
Actions+effects
789.85
58.02
0.027
CoT
685.86
54.37
0.024
ReAct
1028.79
81.555
0.036
ReAct (multi-goal)
3086.37
244.665
0.107
@@
Image: [Uncaptioned image] Naive Prompting: You are navigating the environment below, each elevated object is an obstacle that you have to avoid, Go from point 1 to point 2. Point 1 is located in the top left corner and point 2 is located in the bottom right corner. Your only available actions are (left, right, up, down). Your answer should consist of the predicted sequence solely. Naive Prompting Output: down, down, down, down, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left Remarks: Completely random sequence. Consistent with our finding that naive prompting fails at conceptualizing the task and environment.
@@
Image: [Uncaptioned image] Action and Effects Prompting: You are navigating the environment below, each elevated object is an obstacle that you have to avoid, Go from point 1 to point 2. Your only available actions are (left, right, up, down). Find a path to go from point 1 to point 2. Point 1 is located in the top left corner and point 2 is located in the bottom right corner. Show me the predicted action sequence and what you think would happen with every action Action and Effects Prompting Output: Action: Move down. Outcome: You would move closer to the first obstacle row, avoiding collision. Action: Move down. Outcome: You would be positioned just before the first horizontal obstacle, still avoiding collision. Action: Move right. Outcome: You would navigate towards the gap between the first and second obstacle in the first row. Action: Move down. Outcome: You would now be aligned with the gap between the obstacles, moving closer to the second row. Action: Move right. Outcome: You would enter the second row, moving through the gap between the first and second obstacles. Action: Move down. Outcome: You would be positioned just before the obstacles in the second row. Action: Move right, and continue this action. Outcome: You would pass through the gaps between the obstacles in the subsequent rows, moving down each row closer to point 2. Action: Once aligned with point 2, move right. Outcome: You would reach point 2 located at the bottom right corner of the environment. Remarks: With this prompting method, GPT-4v is able to better reason about the environment and how to avoid obstacles, but the plan generated falls short, as it does not reach the goal.
@@
Image: [Uncaptioned image] Chain of Thought Prompting: You are navigating the environment below, each elevated object is an obstacle that you have to avoid, Go from point 1 to point 2. Point 1 is located in the top left corner and point 2 is located in the bottom right corner. Your only available actions are (left, right, up, down). Provide a sequence of actions to reach point 2 from point 1. Think step-by-step Chain of Thought Prompting Output: Start at Point 1. Move right until the path is obstructed by an obstacle. When an obstacle is directly to the right, move down until you are able to move right again. Continue moving right until you encounter another obstacle. Repeat the process of moving down and then to the right until you can no longer move right because the path is either clear or because you have reached the last column before Point 2. Once in the column adjacent to Point 2, move down until you are aligned horizontally with Point 2. Move right into Point 2. Remarks: The predicted approach outlines a strategy based on environment feedback (e.g. Move right until the path is obstructed by an obstacle.), decomposing the task into segments that require moving in a single direction until an obstacle is encountered; which highlights that when these methods are augmented with a method similar to ReAct to identify when an obstacle has been encountered, it can successfully solve the task. Nevertheless, the solution generated is not optimal as it requires naively reasoning locally to avoid obstacles.

### v3

@@ 1 Introduction
Our experimental results suggest that the GPT-4 model does possess some capability to perform spatial reasoning, when it is prompted effectively. Specifically, the action-and-effect prompting leads to 21.5% improvement in success rate than the naive few-shot prompting, implying that guiding LLMs to reason about the situated knowledge can help them make better immediate decisions. Prompting GPT-4 to reason step-by-step (i.e., CoT) further enhances its performance by 3%, whereas additionally allowing the model to perceive the environment signals (i.e., ReAct) yields the best success rate, 96.1%. However, we also note that this success is mostly limited to relatively “local” reasoning; for example, the ReAct-prompted GPT-4 is shown to plan only a few steps ahead and have to frequently adjust its plan based on the environment signals. In summary, this indicates that even the state-of-the-art GPT-4 still falls short in long-term, temporal reasoning. In contrast, while the fine-tuned LLMs can achieve a close-to-perfect success rate when planning in environments similar to where they were trained (i.e., in-distribution environments), they suffer from the lack of generalization and cannot perform well in environments with different grid sizes or different number of obstacles. We will release our benchmark, along with the source code to controllably synthesize the grid environment, as well as our LLM implementation, to the public for future advancement.  [^1] Our source code and data are available on the following link.
@@ 2.2 Data Generation Process
(Patel & Pavlick, 2022) Understanding None 2D Grid Synthetic No Yes
Ruis et al. (2020) Reasoning Short-horizon Planning 2D Grid Synthetic No No
Shridhar et al. (2020) Reasoning Short-horizon planning Embodied Environment Human-annotated No No
Côté et al. (2018) Reasoning Short-horizon planning Embodied Environment Synthetic No Yes
Shridhar et al. (2021) Reasoning Short-horizon planning Embodied Environment Human-annotated No Yes
Valmeekam et al. (2022) Reasoning Short-horizon planning Blocksworld Synthetic Yes Yes
PPNL (ours) Reasoning Long-horizon Planning 2D Grid Synthetic Yes Yes
@@ 3 Methodologies
ReAct Prompting: Finally, we emulate how an agent can interleavedly perceive environment signals, reason about its plan, and take an action, following the approach of ReAct (Yao et al., 2023). Specifically, we provide two types of signals to the LLM: (1) informing the LLM when it has successfully reached the goal, and (2) reminding the LLM when it is one step ahead of any obstacle, as well as providing its current location. Particularly for the latter type, it allows the LLM to prevent potential failures and correct its ineffective plan. It is important to note that both signals are “local”; that is, the agent will only observe its surroundings and will not be provided with a global view. Therefore, to succeed in the path planning task, it is still required to perform long-term spatial-temporal reasoning. Due to the cost of prompting an LLM iteratively, in experiments, we allow a ReAct LLM to incorporate environment signals and adjust its plan for at most three trials.
@@ 4.1 In-Distribution Experiments
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@ 6-11 Obstacles
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@ 4.2 Out-of-Distribution Experiments
Fine-tuned models can generalize to simpler grids but not to more complex ones. We observed that the fine-tuned LLMs can generalize well to the 5x5 environments. We hypothesize that this is because 5x5 environments are inherently part of the 6x6 environments encountered in training. For the 7x7 environments, however, all of the fine-tuned models failed to perform well. This highlights OOD generalization as an important limitation of fine-tuned models.
@@ 5 Multi-Goal Path Planning Results
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@ 6-11 Obstacles
We notice a trend similar to that of the single-goal setting for OOD generalization. T5-base is still able to perform well on the 5x5 environments; however, it failed catasrophically when faced with larger grids, while also experiencing slighter performance drops on the 6-11 obstacles set.
@@ 7 Related Work
Path and Motion Planning: Many works have assessed whether LLMs can serve as planners (Song et al., 2022; Huang et al., 2022a; Valmeekam et al., 2022). However, many arrive to the conclusion LLMs are unsuitable to be end-to-end planners (Liu et al., 2023; Chen et al., 2023; Xie et al., 2023; Silver et al., 2022). We perform an in-depth analysis to verify this claim. Robot motion and path planning is an exciting application for LLMs that has been explored recently (Ahn et al., 2022; Ding et al., 2023; Chen et al., 2023; Driess et al., 2023; Huang et al., 2022b; Lu et al., 2023; Zhang et al., 2024). Nevertheless, real-world applications of LLMs in the realm of robotics remain limited. Our benchmark can, hence, be particularly useful to the robotic path and motion planning community, as we provide a controlled environment for studying the capability of LLMs to serve as path planners, and their ability to generalize to novel situations, a skill that is crucial in many robotic applications.
@@ A.1 Samples Generation
World Description You are in a $\{N\}$ by $\{N\}$ world.
@@
Enumerating Goals $p\{i\}$ is located at $\{(x_{i},y_{i})\}$
Single goal Go from $\{(x_{init},y_{init})\}$ to $\{(x_{goal},y_{goal})\}$.
Multi Goal Visit the following locations: $\{locations\}$.
Constrained Ordering Visit $\{locations_{a}\}$ before $\{locations_{b}\}$
Initial Location You are at $\{(x_{init},y_{init})\}$
@@ Appendix C Fine-tuned Models’ Performance on Sampled Test Sets
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@ 6-11 Obstacles
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@ Appendix D Additional Analysis
Success ($\uparrow$) Optimal ($\uparrow$) Exact Match ($\uparrow$) Feasible ($\uparrow$) Distance ($\downarrow$) Unreachable Acc ($\uparrow$)
@@
Naive few-shot (5) 418.85 3.91 0.012
Naive few-shot (10) 722.85 3.86 0.022
Naive few-shot (15) 1013.86 3.95 0.031
Actions+effects 789.85 58.02 0.027
CoT 685.86 54.37 0.024
ReAct 1028.79 81.555 0.036
ReAct (multi-goal) 3086.37 244.665 0.107
@@
Planning with GPT-4V(ision): Naive Prompting Image: [Uncaptioned image] Naive Prompting: You are navigating the environment below, each elevated object is an obstacle that you have to avoid, Go from point 1 to point 2. Point 1 is located in the top left corner and point 2 is located in the bottom right corner. Your only available actions are (left, right, up, down). Your answer should consist of the predicted sequence solely. Naive Prompting Output: down, down, down, down, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left Remarks: Completely random sequence. Consistent with our finding that naive prompting fails at conceptualizing the task and environment.
@@
Planning with GPT-4V(ision): Action-Effect Prompting Image: [Uncaptioned image] Action and Effects Prompting: You are navigating the environment below, each elevated object is an obstacle that you have to avoid, Go from point 1 to point 2. Your only available actions are (left, right, up, down). Find a path to go from point 1 to point 2. Point 1 is located in the top left corner and point 2 is located in the bottom right corner. Show me the predicted action sequence and what you think would happen with every action Action and Effects Prompting Output: Action: Move down. Outcome: You would move closer to the first obstacle row, avoiding collision. Action: Move down. Outcome: You would be positioned just before the first horizontal obstacle, still avoiding collision. Action: Move right. Outcome: You would navigate towards the gap between the first and second obstacle in the first row. Action: Move down. Outcome: You would now be aligned with the gap between the obstacles, moving closer to the second row. Action: Move right. Outcome: You would enter the second row, moving through the gap between the first and second obstacles. Action: Move down. Outcome: You would be positioned just before the obstacles in the second row. Action: Move right, and continue this action. Outcome: You would pass through the gaps between the obstacles in the subsequent rows, moving down each row closer to point 2. Action: Once aligned with point 2, move right. Outcome: You would reach point 2 located at the bottom right corner of the environment. Remarks: With this prompting method, GPT-4v is able to better reason about the environment and how to avoid obstacles, but the plan generated falls short, as it does not reach the goal.
@@
Planning with GPT-4V(ision): Chain-of-Thought Prompting Image: [Uncaptioned image] Chain of Thought Prompting: You are navigating the environment below, each elevated object is an obstacle that you have to avoid, Go from point 1 to point 2. Point 1 is located in the top left corner and point 2 is located in the bottom right corner. Your only available actions are (left, right, up, down). Provide a sequence of actions to reach point 2 from point 1. Think step-by-step Chain of Thought Prompting Output: Start at Point 1. Move right until the path is obstructed by an obstacle. When an obstacle is directly to the right, move down until you are able to move right again. Continue moving right until you encounter another obstacle. Repeat the process of moving down and then to the right until you can no longer move right because the path is either clear or because you have reached the last column before Point 2. Once in the column adjacent to Point 2, move down until you are aligned horizontally with Point 2. Move right into Point 2. Remarks: The predicted approach outlines a strategy based on environment feedback (e.g. Move right until the path is obstructed by an obstacle.), decomposing the task into segments that require moving in a single direction until an obstacle is encountered; which highlights that when these methods are augmented with a method similar to ReAct to identify when an obstacle has been encountered, it can successfully solve the task. Nevertheless, the solution generated is not optimal as it requires naively reasoning locally to avoid obstacles.